OpenStudy (anonymous):

7 - 5[(5 + 3)2 - (22 + 10)] ok now this is just plain HARD.... D; can someone show me the steps please?

OpenStudy (anonymous):

calculate the brackets first 7 - 5(8)2 - 32 = 7 - 5*8*2 - 32 next do multiplying =7 - 80 - 32 = -73-32 = - 105

OpenStudy (anonymous):

no, i dont think that is right :c these r the only options;;-383 -343 -93 -243

OpenStudy (anonymous):

a memory aid is useful for doing these type of problens BODMAS - brackets over divide/multiply then add/subtract

OpenStudy (anonymous):

oh - ok i;ll recheck

OpenStudy (anonymous):

yes ii didnt notice the square brackets outside the round ones -

OpenStudy (anonymous):

ok do round brackets first: = 7 - 5[ 8 * 2 - 32] = 7 - 5*(-16) = 7 + 80 = 87

OpenStudy (anonymous):

still not one of your options - another re-check!

OpenStudy (anonymous):

radar - can you check my work plz?

OpenStudy (radar):

O.K

OpenStudy (phi):

I think the problem was typed in wrong. I might not do arithmetic very well, but wolfram can, and it gets 87

OpenStudy (anonymous):

yes

OpenStudy (radar):

87 here also

OpenStudy (anonymous):

right thanks radar - @angie - either problem was typed wrong or the options were

OpenStudy (anonymous):

no. there is a site called "study island" and the answer was 243.

OpenStudy (anonymous):

well there's definitely an inconsistency somewhere - the problem you gave us definitely works out to 87. wolfram does not err - unlike us humans!!